### v1

@@ 1 Introduction
$\langle\nabla F(x^{k}),d^{k}\rangle\leq h\|\nabla F(x^{k})\|^{2},$
@@
where $\mathcal{F}:\mathbf{R}^{n}\to\mathbf{R}^{m}$ is a continuously differentiable function, $E\subset\mathbf{R}^{m}$ is a closed, convex and pointed cone with nonempty interior. Here, the concept of optimality has to be replaced by the concept of Pareto-optimality or efficiency (see Section 2 below).
@@
Very recently, the extensions of the CG methods FR, CD, DY, PRP and HS for vector optimization were first proposed in [29]. Subsequently, the Hager-Zhang CG method [19], the Liu-Storey CG method [18], the Dai-Liao CG method [24] and the spectral CG method [21] for vector optimization were established. It is important to note that in these works, the extensions of the CG method for vector optimization may not always yield descent directions, even when employing the exact or Wolfe-type line search procedure. This phenomenon is prevalent in vector optimization, as illustrated by the examples provided in [19, 21, 24]. To ensure the descent property of search directions in these studies, the redefinition of conjugate parameters is required, or sufficiently accurate line searches are needed, which means that the descent property of search directions depends on the line search method. Therefore, it is valuable to explore CG methods where the generated search direction inherently satisfies the descent property without relying on the line search procedure.
@@
Motivated by the above works, we first propose a three-term PRP-CG method for vector optimization. The main advantage of this method is that the search direction satisfies the sufficient descent property without redefining the conjugate parameter and depending on the line search procedure, while preserving the conjugate property. It is emphasized that this work can be seen as the first generalization of the traditional three-term CG method to vector optimization, which is the main novelty of this paper. In addition, compared with the stronger line search procedures used in [19, 29], we use the generalized Wolfe line search procedure, without the self-adjusting strategy (e.g., [19, 21]), regular restart (e.g., [18]) and convexity assumption (e.g., [24]), establishing the global convergence of the method. The performance of numerical experiments indicates that the proposed scheme has the excellent characteristics of traditional three-term CG methods for scalar optimization.
@@
The rest of this paper is organized as follows. In Section 2, some notations, concepts and preliminary results for vector optimization are presented. In Section 3, a vector extension of the three-term PRP-CG method is proposed, where the search direction naturally satisfies the sufficient descent property independent of any line search procedure. In Section 4, the global convergence of a three-term PRP-CG algorithm for vector optimization is established using the generalized Wolfe line search method. In Section 5, numerical experiments are provided to show the applicability of the scheme. Final remarks and future directions are discussed in Section 6.
@@ 2 Preliminaries
where $\mathcal{F}:\mathbf{R}^{n}\to\mathbf{R}^{m}$ is a continuously differentiable function. We call a point $x^{*}\in\mathbf{R}^{n}$ $E$-Pareto-optimal or $E$-efficient iff there is no $x\in\mathbf{R}^{n}$ such that $\mathcal{F}(x)\preceq_{E}\mathcal{F}(x^{*})$ and $\mathcal{F}(x)\neq\mathcal{F}(x^{*})$. A necessary condition for the local Pareto-optimality of $x\in\mathbf{R}^{n}$ is that
@@
where $J\mathcal{F}(x)$ is the Jacobian of $\mathcal{F}$ at $x$ and Image$(J\mathcal{F}(x))$ is the image of $\mathbf{R}^{n}$ by $J\mathcal{F}(x)$.
A point $x^{*}\in\mathbf{R}^{n}$ is called $E$-critical for $\mathcal{F}$ when it satisfies condition (2.2). Hence, if $x\in\mathbf{R}^{n}$ is not $E$-critical, there exists $y\in\mathbf{R}^{n}$ such that $J\mathcal{F}(x)y\in-\text{int}(E)$, i.e., there exists $\varepsilon>0$ such that
@@
We will denote by $E^{*}$ the positive polar cone of $E$, which is given by
@@
Since $E$ is closed and convex, we have $E=E^{**}$,
@@
In classical optimization (multiobjective optimization), $E=E^{*}=\mathbf{R}_{+}(E=E^{*}=\mathbf{R}_{+}^{m})$, and $V$ can be chosen as $\{1\}$ (a canonical basis of $\mathbf{R}^{m}$).
For the general closed, convex and pointed cone $E$ with nonempty interior, the set
@@
has the properties of $\text{cone}(\text{co}(V))=E^{*}$ and $0\notin$ co$(V)$. In what follows, $V$ will denote exactly this set.
Define $\Psi:\mathbf{R}^{m}\to\mathbf{R}$ and $\Phi:\mathbf{R}^{n}\times\mathbf{R}^{n}\to\mathbf{R}$ by
$\Psi(y):=\sup\{\langle y,v\rangle:v\in V\}$
and
@@
Owing to the compactness of $V$, the function $\Psi$ is well defined, and so is the function $\Phi$. It can be easily checked that
  • $x$ is an $E$-critical point for $\mathcal{F}$ iff $\Phi(x,d)\geq 0$ for all $d\in\mathbf{R}^{n}$;
@@
Since $\Phi(x,\cdot)$ is a real closed convex function, $\mathbf{d}(x)$ exists and is unique. In multiobjective optimization, we choose $E=\mathbf{R}^{m}_{+}$ and $V$ as a canonical basis of $\mathbf{R}^{m}$, then $\mathbf{d}(x)$ can be obtained by solving the following convex quadratic problem:
@@
The following lemma shows that the functions $\mathbf{d}$ and $\Theta$ can be used to characterize the $E$-critical points of problem (2.1).
Lemma 2.2 ([29]).
  1. i
    If $x$ is $E$-critical, then $\mathbf{d}(x)=0$ and $\Theta(x)=0$.
  2. ii
    If $x$ is not $E$-critical, then $\mathbf{d}(x)\neq 0$ is an $E$-descent direction for $\mathcal{F}$ at $x$, $\Theta(x)<0$ and
    $\Phi(x,\mathbf{d}(x))<-\frac{\|\mathbf{d}(x)\|^{2}}{2}<0.$
  3. iii
    The functions $\mathbf{d}$ and $\Theta$ are continuous.
@@
    $\mathcal{F}(x+\alpha d)\preceq\mathcal{F}(x)+\rho\alpha\Phi(x,d)\xi,$
@@
    $\mathcal{F}(x+\alpha d)\preceq\mathcal{F}(x)+\rho\alpha\Phi(x,d)\xi,$ (2.6)
@@
where $10<\rho<\sigma<$ 1 are given constants, $\xi\in E$ is such that for any $v\in V$,
$0<\langle v,\xi\rangle\leq 1$ (2.7)
@@
$\mathcal{F}(x+\alpha d)\preceq\mathcal{F}(x)+\rho\alpha\Phi(x,d)\xi,$ (2.8)
@@ 3 Three term PRP-CG method for vector optimization
In this section, we present a three-term PRP-CG method for vector optimization and describe its characteristics.
@@
which is first proposed in [29] for solving problem (2.1) by Lucambio Pérez and Prudente. Under the descent assumptions, they establish the convergence of the PRP-CG method with the recurrence
@@
In order for CG methods to be well-defined, $d^{k}$ should be a descent direction of $\mathcal{F}$ at $x^{k}$ for all $k\geq 0$, which is equivalent to $\Phi(x^{k},d^{k})<0$ in vector optimization. Some of these methods need the more stringent condition
@@
$d^{k}:=\left\{\begin{array}[]{l}\mathbf{d}(x^{k})\qquad\text{for}\ k=0,\\ \mathbf{d}(x^{k})+\beta^{PRP}_{k}d^{k-1}+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\qquad\text{for}\ k\geq 1.\end{array}\right.$ (3.4)
@@
If $x^{k}$ is not $E$-critical and the search direction $d^{k}$ is defined as in (3.4), then $d^{k}$ satisfies the sufficient descent condition (3.3) with $h=1$.
@@
Since $x^{k}$ is not $E$-critical, we obtain $\mathbf{d}(x^{k})\neq 0$ by Lemma 2.2(ii). The condition (3.3) trivially holds for $d^{0}=\mathbf{d}(x^{0})$. For any $k\geq 1$, it can be deduced from (3.1) that
$\Phi(x^{k},d^{k})=\Phi\bigg{(}x^{k},\mathbf{d}(x^{k})+\beta^{PRP}_{k}d^{k-1}+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{)}.$ (3.5)
We now consider the following cases:
Case 1: $\beta^{PRP}_{k},\Phi(x^{k},d^{k-1})\geq 0$. By the definition of $\Phi$ and the fact $\Phi(x^{k},\mathbf{d}(x^{k}))<0$ obtained from Lemma 2.2(ii), we have
$\Phi(x^{k},-\mathbf{d}(x^{k}))$ $=\sup\{\langle-J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z\rangle:z\in V\}$ (3.6)
$\geq\langle-J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z_{0}\rangle$
$=-\langle J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z_{0}\rangle$
$=-\Phi(x^{k},\mathbf{d}(x^{k}))>0,$
where $z_{0}\in V$ is such that
$\langle J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z_{0}\rangle=\sup\{\langle J\mathcal{F}(x^{k})\mathbf{d}(x^{k}),z\rangle:z\in V\}=\Phi(x^{k},\mathbf{d}(x^{k})).$
Combining (3.5), Lemma 2.1(i) and (3.6), we have
$\Phi(x^{k},d^{k})$ $\leq\Phi(x^{k},\mathbf{d}(x^{k}))+\beta^{PRP}_{k}\Phi(x^{k},d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k}))+\beta^{PRP}_{k}\Phi(x^{k},d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{-\Phi(x^{k},\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
@@
Case 2: $\beta^{PRP}_{k}\geq 0$ and $\Phi(x^{k},d^{k-1})<0$. It follows from (3.5) and Lemma 2.1(i) that
$\Phi(x^{k},d^{k})$ $\leq\Phi(x^{k},\mathbf{d}(x^{k}))+\beta^{PRP}_{k}\Phi(x^{k},d^{k-1})+\beta^{PRP}_{k}\frac{-\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},-\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k})).$
Case 3: $\beta^{PRP}_{k},\Phi(x^{k},d^{k-1})<0$. Using similar arguments in (3.6), we infer that
$\Phi(x^{k},-d^{k-1})>-\Phi(x^{k},d^{k-1})>0,$
which, combined with (3.5), Lemma 2.1(i) and (3.6), implies that
$\Phi(x^{k},d^{k})$ $\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},d^{k-1})}{-\Phi(x^{k},\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},-d^{k-1})}{\Phi(x^{k},\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k})).$
Case 4: $\beta^{PRP}_{k}<0,\Phi(x^{k},d^{k-1})\geq 0$. Employing similar lines in (3.6), we have
$0>\Phi(x^{k},-d^{k-1})>-\Phi(x^{k},d^{k-1}),$
which, along with (3.5) and Lemma 2.1(i), yields that
$\Phi(x^{k},d^{k})$ $\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{-\Phi(x^{k},d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},-\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k}))-\beta^{PRP}_{k}\Phi(x^{k},-d^{k-1})+\beta^{PRP}_{k}\frac{\Phi(x^{k},-d^{k-1})}{\Phi(x^{k},-\mathbf{d}(x^{k}))}\Phi(x^{k},-\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k})).$
Combining the above cases, we obtain $\Phi(x^{k},d^{k})\leq\Phi(x^{k},\mathbf{d}(x^{k}))$, which establishes the theorem.

### v2

@@ 1 Introduction
$\langle\nabla F(x^{k}),d^{k}\rangle\leq h\|\nabla F(x^{k})\|^{2}$
@@
where $\mathcal{F}:\mathbf{R}^{n}\to\mathbf{R}^{m}$ is a continuously differentiable function, $E\subset\mathbf{R}^{m}$ is a closed, convex and pointed cone with nonempty interior. Here, the concept of optimality has to be replaced by the concept of Pareto-optimality (see Section 2 below). Moreover, when $E=\mathbf{R}^{m}_{+}$, this problem can be reduced to a multi-objective optimization problem that finds acceptable trade-offs between the objectives. For the work on multi-objective optimization, we refer to [5, 16, 18, 19, 31, 33, 40] and the references therein.
@@
Very recently, the extensions of the CG methods FR, CD, DY, PRP and HS for vector optimization were first proposed in [36]. Subsequently, the Hager-Zhang CG method [23], the Liu-Storey CG method [22], the Dai-Liao CG method [28] and the spectral CG method [25] for vector optimization were established. It is important to note that in these works, the extensions of the CG method for vector optimization may not always yield descent directions, even when employing the exact or Wolfe-type line search procedure. This phenomenon is prevalent in vector optimization, as illustrated by the examples provided in [23, 25, 28]. To ensure the descent property of search directions in these studies, the redefinition of conjugate parameters is required, or sufficiently accurate line searches are needed, which means that the descent property of search directions depends on the line search method. Therefore, it is valuable to explore CG methods where the generated search direction inherently satisfies the descent property without relying on the line search procedure. The variations of the PRP-CG method for vector optimization have been discussed in [6, 29] recently. More research on CG methods for vector optimization was considered in [7, 30, 45, 46]. However, as far as we know, neither the three-term PRP-CG method nor the generalized three-term CG method has been extended to vector optimization, which is the motivation for the current study.
@@
Motivated by the above works, we propose for the first time a three-term PRP-CG method for vector optimization. The main advantage of this method is that the search direction satisfies the sufficient descent property without redefining the conjugate parameter and depending on the line search procedure, while preserving the conjugate property. It is emphasized that this work can be seen as the first generalization of the traditional three-term CG method to vector optimization, which is the main novelty of this paper. In addition, compared with the stronger line search procedures used in [23, 36], we use the generalized Wolfe line search procedure, without the self-adjusting strategy (e.g., [23, 25]), regular restart (e.g., [22]) and convexity assumption (e.g., [28]), establishing the global convergence of the method. The performance of numerical experiments indicates that the proposed scheme has the excellent characteristics of traditional three-term CG methods for scalar optimization.
The main contributions of this paper are as follows:
  • The three-term PRP-CG method has, for the first time, been extended to vector optimization. This is also the first application of three-term CG methods in the field of vector optimization. The extension retains the excellent property of three-term CG methods in scalar optimization, i.e., the search direction has the sufficient descent property without any line search method. Furthermore, this extension addresses the gap of the three-term CG method for vector optimization.
  • By introducing the generalized Wolfe line search method, the proposed algorithm eliminates the reliance on the strong line search procedure, self-adjusting strategy, regular restart and convexity assumption used in common methods. Under more relaxed conditions, the global convergence of the algorithm is established.
@@
The rest of this paper is organized as follows: In Section 2, some notations, concepts and preliminary results for vector optimization are presented. In Section 3, a vector extension of the three-term PRP-CG method is proposed, where the search direction naturally satisfies the sufficient descent property independent of any line search procedure. In Section 4, the global convergence of a three-term PRP-CG algorithm for vector optimization is established using the generalized Wolfe line search method. In Section 5, numerical experiments are provided to show the applicability of the scheme. Final remarks and future directions are discussed in Section 6.
@@ 2 Preliminaries
where $\mathcal{F}:\mathbf{R}^{n}\to\mathbf{R}^{m}$ is a continuously differentiable function. We call a point $x^{*}\in\mathbf{R}^{n}$ $E$-Pareto-optimal (resp. weak $E$-Pareto-optimal) iff there is no $x\in\mathbf{R}^{n}$ such that $\mathcal{F}(x)\preceq_{E}\mathcal{F}(x^{*})$ (resp. $\mathcal{F}(x)\prec_{E}\mathcal{F}(x^{*})$) and $\mathcal{F}(x)\neq\mathcal{F}(x^{*})$.
A necessary condition for the local (weak) $E$-Pareto-optimality of $x\in\mathbf{R}^{n}$ is that
@@
where $J\mathcal{F}(x)$ denotes the Jacobian of $\mathcal{F}$ at $x$ and Image$(J\mathcal{F}(x))$ represents the image of $\mathbf{R}^{n}$ by $J\mathcal{F}(x)$. A point $x^{*}\in\mathbf{R}^{n}$ is called $E$-Pareto-critical for $\mathcal{F}$ if it satisfies condition (2.2). Note that if $x\in\mathbf{R}^{n}$ is not $E$-Pareto-critical, there exists $y\in\mathbf{R}^{n}$ such that $J\mathcal{F}(x)y\in-\text{int}(E)$, i.e., there exists $\varepsilon>0$ such that
@@
We denote the positive polar cone of $E$ as
@@
Since $E$ is closed and convex, we use the definition of the positive polar cone to obtain $E=E^{**}$ ,
@@
For a general closed, convex and pointed cone $E$ with nonempty interior, the set
@@
has the properties of $\text{cone}(\text{co}(V))=E^{*}$ and $0\notin V$. In what follows, $V$ will denote exactly this set.
Remark 2.1.
In multiobjective optimization, $E=E^{*}=\mathbf{R}_{+}^{m}$ and $V$ can be chosen as a canonical basis of $\mathbf{R}^{m}$.
Define $\Phi:\mathbf{R}^{n}\times\mathbf{R}^{n}\to\mathbf{R}$ by
@@
Owing to the compactness of $V$, the function $\Phi$ is well defined. It can be easily checked that
  • $x$ is $E$-Pareto-critical for $\mathcal{F}$ iff $\Phi(x,d)\geq 0$ for all $d\in\mathbf{R}^{n}$;
@@
Since $\Phi(x,\cdot)$ is a real closed convex function, $\mathbf{d}(x)$ exists and is unique. The following lemma shows that the functions $\mathbf{d}$ and $\Theta$ can be used to describe $E$-Pareto-critical points.
Lemma 2.3 ([15]).
  1. (i)
    If $x$ is $E$-Pareto-critical, then $\mathbf{d}(x)=0$ and $\Theta(x)=0$.
  2. (ii)
    If $x$ is not $E$-Pareto-critical, then $\mathbf{d}(x)\neq 0$ is an $E$-descent direction for $\mathcal{F}$ at $x$, $\Theta(x)<0$ and
    $\Phi(x,\mathbf{d}(x))<-\frac{\|\mathbf{d}(x)\|^{2}}{2}<0.$
  3. (iii)
    The functions $\mathbf{d}$ and $\Theta$ are continuous.
Remark 2.4.
In multiobjective optimization, we choose $E=\mathbf{R}^{m}_{+}$ and $V$ as a canonical basis of $\mathbf{R}^{m}$, then $\mathbf{d}(x)$ can be obtained by solving the convex quadratic problem:
@@
    $\mathcal{F}(x+\alpha d)\preceq_{E}\mathcal{F}(x)+\rho\alpha\Phi(x,d)\xi,$
@@
    $\mathcal{F}(x+\alpha d)\preceq_{E}\mathcal{F}(x)+\rho\alpha\Phi(x,d)\xi,$ (2.6)
@@
where $10<\rho<\sigma<$ 1 are given constants, $\xi\in E$ is such that
$0<\langle\xi,v\rangle\leq 1\qquad\forall v\in V.$ (2.7)
@@
$\mathcal{F}(x+\alpha d)\preceq_{E}\mathcal{F}(x)+\rho\alpha\Phi(x,d)\xi,$ (2.8)
@@ 3 Three term PRP-CG method for vector optimization
In this part, we present a three-term PRP-CG method for vector optimization and describe its characteristics.
@@
which is first proposed in [36] for solving the unconstrained vector optimization problem (2.1) by Lucambio Pérez and Prudente. Under the descent assumptions, they establish the convergence of the PRP-CG method with the recurrence
@@
In order for CG methods to be well defined, $d^{k}$ should be a descent direction of $\mathcal{F}$ at $x^{k}$ for all $k\geq 0$, which is equivalent to $\Phi(x^{k},d^{k})<0$ in vector optimization. Some of these methods need the more stringent condition
@@
$d^{k}:=\left\{\begin{array}[]{l}\mathbf{d}(x^{k})\qquad\text{for}\ k=0,\\ \mathbf{d}(x^{k})+\beta^{PRP+}_{k}d^{k-1}-\beta^{PRP+}_{k}\frac{|\Phi(x^{k},d^{k-1})|}{\Phi(x^{k},\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\qquad\text{for}\ k\geq 1.\end{array}\right.$ (3.4)
@@
If $x^{k}$ is not $E$-Pareto-critical and the search direction $d^{k}$ is defined as in (3.4), then $d^{k}$ satisfies the sufficient descent condition (3.3) with $h=1$.
@@
Since $x^{k}$ is not $E$-Pareto-critical, we obtain $\mathbf{d}(x^{k})\neq 0$ by Lemma 2.3(ii). The condition (3.3) trivially holds for $d^{0}=\mathbf{d}(x^{0})$. For any $k\geq 1$, it can be deduced from (3.1) that
$\Phi(x^{k},d^{k})=\Phi\bigg{(}x^{k},\mathbf{d}(x^{k})+\beta^{PRP+}_{k}d^{k-1}-\beta^{PRP+}_{k}\frac{|\Phi(x^{k},d^{k-1})|}{\Phi(x^{k},\mathbf{d}(x^{k}))}\mathbf{d}(x^{k})\bigg{)}.$ (3.5)
By Lemma 2.2(i) and the fact $\Phi(x^{k},\mathbf{d}(x^{k}))<0$ obtained from Lemma 2.3(ii), we can infer from (3.5) that
$\Phi(x^{k},d^{k})$ $\leq\Phi(x^{k},\mathbf{d}(x^{k}))+\beta^{PRP+}_{k}\Phi(x^{k},d^{k-1})-\beta^{PRP+}_{k}\frac{|\Phi(x^{k},d^{k-1})|}{\Phi(x^{k},\mathbf{d}(x^{k}))}\Phi(x^{k},\mathbf{d}(x^{k}))$
$\leq\Phi(x^{k},\mathbf{d}(x^{k}))+\beta^{PRP+}_{k}\Phi(x^{k},d^{k-1})-\beta^{PRP+}_{k}|\Phi(x^{k},d^{k-1})|$
@@
Combining these, we obtain $\Phi(x^{k},d^{k})\leq\Phi(x^{k},\mathbf{d}(x^{k}))$, which establishes the theorem.